# A solution to Banach conjecture

### v4

@@ 3. Proof of the main theorem
Theorem 3.1 follows from the classical characterization by Busemann of ellipsoids (see [14, Theorem 3.1]). The convexity of the body $K$ results from the fact that if we have two points in the body, we can examine the plane they span (with the origin). Since the planar section is an ellipsoid by assumption, the entire line segment between $x$ and $y$ is contained in $K$.
@@
First, we introduce two functions $\Lambda:\mathbb{S}^{n-1}\to\mathbb{R}_{\geq 0}$ and $\Gamma:G(n,n-2)\to\mathbb{R}_{\geq 0}$ defined by
$\Lambda(\theta):=\sup_{\xi,\eta\in\mathbb{S}^{n-1}\cap\theta^{\perp}}\Big\{\Big|\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}-\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}\Big|\Big\}$
and
@@
It is clear that
$\Gamma(H)=\sup_{\theta\in\mathbb{S}^{n-1}\cap H}\Lambda(\theta)$
and Condition (1) is equivalent to $\Lambda(\theta)=0$ for all $\theta\in\mathbb{S}^{n-1}$ or $\Gamma(H)=0$ for all $H\in G(n,n-2)$.
Moreover, $\Lambda,\Gamma$ is continuous because $\Phi(\xi)$ is continuous and $J(K\cap\xi_{0}^{\perp})$ is invariant under $G_{K\cap\xi_{0}^{\perp}}$. Furthermore, $\Gamma(H_{1})=\Gamma(H_{2})$ whenever there exists a linear transformation between $K\cap H_{1}$ and $K\cap H_{2}$, as implied by Lemma 3.1.
Now we partition G(n,n2)G(n,n-2) into two subsets, $A_{1}$ and $A_{2}$. The set $A_{1}$ consists of all $H$ such that, for every $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$,
$\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap H)=\phi_{\xi}\circ g\circ\phi^{-1}_{\xi}(K\cap H)$
for some $g\in G_{K\cap\xi_{0}^{\perp}}$. The set $A_{2}$ consists of all $H\in G(n-2)$ for which there exists a pair of vectors $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$ such that
$\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap H)\neq\phi_{\xi}\circ g\circ\phi^{-1}_{\xi}(K\cap H)$
@@
We once more partition G(n,n2)G(n,n-2) into two subsets, $B_{1}$ and $B_{2}$. Define
Our first goal is to show that if $B_{2}\neq\emptyset$, $B_{2}=G(n,n-2)$.
@@
$\phi_{\eta}(\phi^{-1}_{\xi}(K\cap H)\cap J(K\cap\xi_{0}^{\perp}))=\phi_{\xi}\circ g(\phi^{-1}_{\xi}(K\cap H)\cap J(K\cap\xi_{0}^{\perp})),$
where $g(J(K\cap\xi_{0}^{\perp}))=J(K\cap\xi_{0}^{\perp})$, and thus, we have
$\phi_{\eta}(\phi^{-1}_{\xi}(H)\cap J(K\cap\xi_{0}^{\perp}))=\phi_{\xi}(\phi^{-1}_{\xi}(H)\cap J(K\cap\xi_{0}^{\perp})),$
@@
$\phi_{\eta}\circ\phi^{-1}_{\xi}(H)\cap\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))=H\cap\phi_{\xi}(J(K\cap\xi_{0}^{\perp})),$
and thus,
$\phi_{\xi}\circ g\circ\phi^{-1}_{\xi}(H)\cap\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))=H\cap\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$
@@
$\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))\cap H=\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))\cap H.$
Therefore, we obtain
@@
$\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap H)=\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})=K\cap\eta^{\perp}\cap\phi_{\eta}\circ\phi^{-1}_{\xi}(\eta^{\perp}).$
Moreover, since $H\in B_{2}$, there is some $\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ such that
$\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})\neq\phi_{\xi}\circ g\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})$
for every $g\in G_{K\cap\xi_{0}^{\perp}}$. In particular, this implies $\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})\neq K\cap\xi^{\perp}\cap\eta^{\perp}$, and by the continuity of $\Phi$, there is a path joining $\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})$ to $K\cap\xi^{\perp}\cap\eta^{\perp}$ by $K\cap\zeta^{\perp}\cap\phi_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp})$ with $\eta$ lying on the arc between $\xi$ and $\eta$.
Thus, as $\zeta$ moves past $\xi$, there are two possibilities: either $K\cap\eta^{\perp}\cap\phi_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp})$ returns to $K\cap\xi_{0}^{\perp}\cap\zeta^{\perp}$ exactly $m$ times, giving rise to an $m$-circle, or $K\cap\eta^{\perp}\cap\phi_{\zeta}\circ\phi^{-1}_{\xi}(\eta^{\perp})$ arrives at $K\cap\xi_{0}^{\perp}\cap\zeta^{\perp}$, meaning that $\lim_{\zeta\to-\xi}\phi_{\zeta}\circ\phi^{-1}_{\xi}(\theta)=-\theta$ for $\theta\in\xi^{\perp}$.
@@
$\mathcal{L}(H):=\{\zeta^{\perp}\cap\phi_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp}):\zeta,\xi\in\mathbb{S}^{n-1}\cap H^{\perp}\}.$
Consequently, $\mathcal{L}(H)\cap\mathbb{S}^{n-1}$ is either a $2m$-circle or a single large circle passing through the antipodal point. Moreover, for any $H_{1},H_{2}\in\mathcal{L}(H)\cap G(n,n-2)$, the sections $K\cap H_{1}$ and $K\cap H_{2}$ are linearly dependent.
We now assume that $B_{2}$ is nonempty, which implies that the maximum value of $\Lambda(\theta)$, denoted by $M$, is nonzero. By the definition of $\Gamma$, for every $H\in G(n,n-2)$ containing $\theta$, we have $\Gamma(H)=M$. Clearly, such an $H$ lies in $B_{2}$, and thus also in $A_{2}$. Hence, by the preceding argument and the fact that $\Gamma$ is invariant under linear transformations, the level set
@@
We now show that if $B_{2}\neq\emptyset$, then in fact $B_{2}=G(n,n-2)$, which implies $A_{2}=G(n,n-2)$. By repeating the preceding argument for $\mathcal{L}(H)$ with any $H\in G(n,n-2)$, we conclude that all (n2)(n-2)-dimensional sections of $K$ are linearly dependent. Consequently, by Gromov’s theorem [7], $K$ must be an ellipsoid. ∎

### v5

@@ 3. Proof of the main theorem
Theorem 3.1 follows from the classical characterization by Busemann of ellipsoids (see [15, Theorem 3.1]). The convexity of the body $K$ results from the fact that if we have two points in the body, we can examine the plane they span (with the origin). Since the planar section is an ellipsoid by assumption, the entire line segment between $x$ and $y$ is contained in $K$.
@@
First, we introduce a function $\Gamma:G(n,n-2)\to\mathbb{R}_{\geq 0}$ defined by
@@
It is clear that Condition (1) is equivalent to $\Gamma(H)=0$ for all $H\in G(n,n-2)$. And hence, We once more partition G(n,n2)G(n,n-2) into two subsets, $B_{1}$ and $B_{2}$. Define
Our goal is to show that if $B_{2}\neq\emptyset$, $B_{2}=G(n,n-2)$.
Moreover, $\Gamma$ is continuous because $\Phi(\xi)$ is continuous and $J(K\cap\xi_{0}^{\perp})$ is invariant under $G_{K\cap\xi_{0}^{\perp}}$. Furthermore, $\Gamma(H_{1})=\Gamma(H_{2})$ whenever there exists a linear transformation between $K\cap H_{1}$ and $K\cap H_{2}$, as implied by Lemma 3.1. Now we need to deal with the linear transformation inside G(n,n2)G(n,n-2).
Observe that, for any $H\in G(n,n-2)$ and any $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$, there exist two linear maps $\phi^{-1}_{\xi},\phi^{-1}_{\eta}$ that send $H$ back to $\xi_{0}^{\perp}$. Depending on the relative positions of $\phi^{-1}_{\xi}(H)$ and $\phi^{-1}_{\eta}(H)$, we divide G(n,n2)G(n,n-2) into two subsets, denoted $A_{1}$ and $A_{2}$.
The set $A_{1}$ consists of all $H$ such that, for every $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$,
$\phi^{-1}_{\eta}(K\cap H)=g\circ\phi^{-1}_{\xi}(K\cap H)$
for some $g\in G_{K\cap\xi_{0}^{\perp}}$.
The set $A_{2}$ consists of all $H\in G(n-2)$ for which there exists a pair of vectors $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$ such that
$\phi^{-1}_{\eta}(K\cap H)\neq g\circ\phi^{-1}_{\xi}(K\cap H)$
@@
$\phi^{-1}_{\eta}(K\cap H)\cap J(K\cap\xi_{0}^{\perp})=g\circ\phi^{-1}_{\xi}(K\cap H)\cap J(K\cap\xi_{0}^{\perp}),$
@@
$\phi^{-1}_{\eta}(K\cap H)\cap J(K\cap\xi_{0}^{\perp})=g(\phi^{-1}_{\xi}(K\cap H)\cap g^{-1}(J(K\cap\xi_{0}^{\perp}))),$
where $g^{-1}(J(K\cap\xi_{0}^{\perp}))=J(K\cap\xi_{0}^{\perp})$, and thus, we have
$\phi^{-1}_{\eta}(K\cap H)\cap J(K\cap\xi_{0}^{\perp})=g(\phi^{-1}_{\xi}(K\cap H)\cap J(K\cap\xi_{0}^{\perp})),$
and hence,
$\phi^{-1}_{\eta}(K\cap H\cap\phi_{\eta}(J(K\cap\xi_{0}^{\perp})))=g(\phi^{-1}_{\xi}(K\cap H\cap\phi_{\xi}(J(K\cap\xi_{0}^{\perp})))),$
@@
$\phi^{-1}_{\eta}(\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))\cap H)=g\circ\phi^{-1}_{\xi}(\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))\cap H).$
Therefore, as a result of $\phi^{-1}_{\eta}(K\cap H)=g\circ\phi^{-1}_{\xi}(K\cap H)$, we obtain
$\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))\cap H=\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))\cap H$
which means
@@
$\phi_{\eta}\circ g\circ\phi^{-1}_{\xi}(K\cap H)=\phi_{\eta}\circ g\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})=K\cap\eta^{\perp}\cap\phi_{\eta}\circ g\circ\phi^{-1}_{\xi}(\eta^{\perp}).$
Moreover, since $H\in A_{2}$, there is some $\eta,\xi\in\mathbb{S}^{n-1}\cap H^{\perp}$ such that
$\phi^{-1}_{\eta}(K\cap\xi^{\perp}\cap\eta^{\perp})\neq g\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})$
for every $g\in G_{K\cap\xi_{0}^{\perp}}$. In particular, we have $\phi_{\eta}\circ g\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})\neq K\cap\xi^{\perp}\cap\eta^{\perp}$, and by the continuity of $\Phi$, there is a path joining $\phi_{\eta}\circ g_{\eta}\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})$ to $K\cap\xi^{\perp}\cap\eta^{\perp}$ by $K\cap\zeta^{\perp}\cap\phi_{\zeta}\circ g_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp})$ with $\zeta$ lying on the arc between $\xi$ and $\eta$.
Thus, as $\zeta$ moves past $\xi$, there are two possibilities: either $K\cap\eta^{\perp}\cap\phi_{\zeta}\circ g_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp})$ returns to $K\cap\xi^{\perp}\cap\zeta^{\perp}$ exactly $m$ times, giving rise to an $m$-circle, or $K\cap\eta^{\perp}\cap\phi_{\zeta}\circ g_{\zeta}\circ\phi^{-1}_{\xi}(\eta^{\perp})$ arrives at $K\cap\xi^{\perp}\cap\zeta^{\perp}$, meaning that $\lim_{\zeta\to-\xi}\phi_{\zeta}\circ g_{\zeta}\circ\phi^{-1}_{\xi}(\theta)=-\theta$ for $\theta\in\xi^{\perp}$.
@@
$\mathcal{L}(H):=\{\zeta^{\perp}\cap\phi_{\zeta}\circ g_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp}):\zeta,\xi\in\mathbb{S}^{n-1}\cap H^{\perp}\}.$
Consequently, $\mathcal{L}(H)\cap\mathbb{S}^{n-1}$ contains either a $2m$-loop or a single large loop passing through the antipodal point. Moreover, for any $H_{1},H_{2}\in\mathcal{L}(H)\cap G(n,n-2)$, the sections $K\cap H_{1}$ and $K\cap H_{2}$ are linearly dependent.
We now assume that $B_{2}$ is nonempty, which implies that the maximum value of $\Gamma(H)$, denoted by $M$, is nonzero. Hence, by the preceding argument and the fact that $\Gamma$ is invariant under linear transformations, the level set
@@
We now show that if $B_{2}\neq\emptyset$, then in fact $B_{2}=G(n,n-2)$, which implies $A_{2}=G(n,n-2)$. In other words, for any $H\in G(n,n-2)$, the set $\mathcal{H}$ either contains an (n1)(n-1)-dimensional manifold with antipodal points, or it contains a subset of the form $H\times\infty$, such that for all $H_{1},H_{2}\in\mathcal{L}(H)\cap G(n,n-2)$, the sections $K\cap H_{1}$ and $K\cap H_{2}$ are linearly dependent.
We now introduce another family $\mathcal{F}(H)$ defined by
$\mathcal{F}(H):=\{\tilde{H}\in G(n,n-2):K\cap\tilde{H}\ \text{and}\ K\cap H\ \text{are linearly dependent}\}.$
Observe that $\mathcal{L}(H)\subseteq\mathcal{F}(H)$. We then assert that there can be only finitely many pairwise disjoint sets $\{H_{i}\in G(n,n-2)\}_{i=1}^{m}$, each containing a subset of the form $H\times\infty$, such that
$\mathcal{F}(H_{i})\cap\mathcal{F}(H_{j})=\emptyset$
for $i\neq j$. Suppose, to the contrary, that there are infinitely many pairwise disjoint sets $\{H_{i}\}_{i=1}^{\infty}$. Then, for each $i$, the set $H_{i+1}\times\infty$ is contained in $C_{i}$, one of the connected components of the complement determined by $H_{i}\times\infty$. By the compactness of G(n,n2)G(n,n-2), both sequences $\{C_{i}\}$ and $\{H_{i}\}$ converge to some limit $\bar{H}$. However, for all sufficiently large $i\geq I$, we have
[…]
which implies
$\mathcal{L}(\bar{H})\cap\mathcal{L}(H_{i})\neq\emptyset$
for all $i\geq I$. This yields a contradiction, since it shows that there must exist a linear transformation relating the $H_{i}$ for all $i\geq I$.
We now choose $H_{0}$ such that $\mathcal{L}(H_{0})$ contains an (n1)(n-1)-dimensional manifold with antipodal points. For any other $H$ for which $\mathcal{L}(H)$ also contains an (n1)(n-1)-dimensional manifold with antipodal points, we obtain
$\mathcal{L}(H_{0})\cap\mathcal{L}(H)\neq\emptyset,$
and consequently,
$\mathcal{L}(H)\subseteq\mathcal{F}(H_{0}).$
If for some $i\in[1,m]$ we have
$\mathcal{F}(H_{0})\cap\mathcal{F}(H_{i})\neq\emptyset,$
then it follows that
$\mathcal{F}(H_{0})=\mathcal{F}(H_{i}).$
Otherwise, for every $i\in[1,m]$,
$\mathcal{F}(H_{0})\cap\mathcal{F}(H_{i})=\emptyset.$
Consequently, G(n,n2)G(n,n-2) is partitioned into $m$ or $m+1$ pairwise disjoint closed sets, and thus, by the connectedness of G(n,n2)G(n,n-2), we conclude that all (n2)(n-2)-dimensional sections of $K$ are linearly dependent. Consequently, by Gromov’s theorem [7], $K$ must be an ellipsoid. ∎